• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name LeRose

    Log entry time 16:00:41 on March23,2005

    Entry number 142257

    keyword=Day Shift Summary

    At start of shift the experiment is down due to an overheating problem in the left Q2 power supply. Techs and Bill Vulcan are down in the Hall trouble shooting.
    At around noon we get a positive diagnosis on the LQ2 power supply: Bad SCR in the power supply, estimate 4 hours to replace.
    At shift's end they are still working on the LQ2 power supply.